OpenStudy (anonymous):

A jet ski company charges a flat fee of $27.00 plus $2.00 per hour to rent a jet ski. Another company

OpenStudy (anonymous):

continue?

OpenStudy (anonymous):

A jet ski company charges a flat fee of $27.00 plus $2.00 per hour to rent a jet ski. Another company charges a fee of $16.50 plus $3.75 per hour to rent the same jet ski.

OpenStudy (anonymous):

Whats the question

OpenStudy (anonymous):

It just says solve the system of equations.

OpenStudy (ddcamp):

Make an equation for each of the Jet ski companies: Cost = base rate + (hourly rate × hours)

OpenStudy (anonymous):

Would the cost be y? I wrote 27+2x=y and 16.5+3.75x=y

OpenStudy (anonymous):

You first need to find two equations. One for each company. Then with your equations, input them into y1 and y2 in your calc. Graph them and find the intersection point.

OpenStudy (anonymous):

and they cost the same at 4 hours

OpenStudy (anonymous):

Thanks! It doesn't say I need to graph it or anything. But would I solve by elimination or substation? Do I put them equal to each other and solve it that way? I have to show my work.

OpenStudy (anonymous):

substitution*

OpenStudy (ddcamp):

You're finding when the cost of both of them is the same. Since cost is y, and you're wanting the cost to be equal, you can set the equations equal to each other: 27+2x=y and 16.5+3.75x=y 27+2x = 16.5 + 3.75x 10.5 = 1.75x 6 = x